Fire hoses used in major structure fires have inside diameters of 6.0 cm. Suppose such a hose carries a flow of 40.0 L/s starting at a gauge pressure of 2.2×106 N/m2. The hose goes 10.0 m up a ladder to a nozzle having an inside diameter of 2.8 cm.
a.   What is the pressure in the nozzle?
b.   What is the speed of water coming out of the nozzle?
Answer #1

a)

V = Volume flow rate = 40 L/s = 40 x 10-3 m3/s

dn = diameter of nozzle = 2.8 cm = 0.028 m

An = area of cross-section of the nozzle = (0.25) \pi dn2 = (0.25) (3.14) (0.028)2 = 0.00061544 m2

vn = speed of water coming out of nozzle = ?

Volume flow rate is given as

V = An vn

(40 x 10-3) = (0.00061544) vn

vn = 65 m/s

dh = diameter of hose = 6 cm = 0.06 m

Ah = area of cross-section of the hose = (0.25) \pi dh2 = (0.25) (3.14) (0.06)2 = 0.002826 m2

vh = speed of water coming out of hose = ?

Volume flow rate is given as

V = Ah vh

(40 x 10-3) = (0.002826) vh

vh = 14.2 m/s

Ph = gauge pressure at hose = 2.2 x 106 Pa

Pn = gauge pressure at nozzle

h1 = 0

h2 = 10 m

Using Bernoulli's theorem

Ph + \rho g h1 = Pn + \rho g h2

2.2 x 106 + (1000) (9.8) (0) = Pn + (1000) (9.8) (10)

Pn = 2.102 x 106 Pa